write an aquation in point-rope form of the line having a y-intercept 7 and passing through (1, 8)
Answer by mananth(16946)

y= mx+b is the equation in slope intercept form
point(1,8) and y intercept =7
plug the values
8=1*m+7
8-7=m
m=1
m=1,b=7
so the equation is y=x+7